OpenStudy (anonymous):

A truck contains 150 small packages, some weighing 1kg each and some weighing 2kg each. How many packages weighing 2kg each are in the truck if the total weigh of all the packages is 264 kg?

OpenStudy (anonymous):

The answer choices are 36, 52, 88, 124, or 114.

OpenStudy (anonymous):

x+y = 150 1x+2y = 264 (solve simultaneously) y=114

OpenStudy (anonymous):

How do you solve 1x+2y=264 ?

OpenStudy (anonymous):

1x + 2y = 264 -(x+y = 150) *to eliminate variable x --------------- y= 114

OpenStudy (anonymous):

Ohh! ok thank you.